OpenStudy (errinss):

What is the value of y? 3y^2 − 6 = 42

OpenStudy (cheez):

lol I put the answer in on the last one. But ill do it again

OpenStudy (errinss):

my computer froze

OpenStudy (cheez):

no prob. Add the six, as I taught you before. Then, divide by 3 and make both sides equal. Then sqrt. And boom - answer.

OpenStudy (errinss):

yup

OpenStudy (cheez):

Great! there you have it! It seems that you should now know how to do this on your own. i'll make up one and I want you to use what I've taught you to figure it out. 3x^2 = 192. AND GO!

OpenStudy (errinss):

x^2=64 so x would be 8

OpenStudy (cheez):

YES! FANTASTIC! You got this! By George I think ye've got it!

OpenStudy (errinss):

i divided 3 by both sides and then square rooted 64 which is 8 because 8*8 is 64

OpenStudy (cheez):

That's how you do it! Great Job, my work here is done.

OpenStudy (errinss):

your username makes me smile, i had a cat named Cheeze

pooja195 (pooja195):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @errinss x^2=64 so x would be 8 \(\color{#0cbb34}{\text{End of Quote}}\) Rememeber that there would be 2 answers, a negative one and a positive one.

OpenStudy (cheez):

Lol! All my usernames are Cheez or Cheezywizard5292 because of my IRL name, its a type of cheese. And I like wizards. and 5292.

OpenStudy (cheez):

Pooja is right, because -8^2 is the exact same as 8^2 because the ^2 cancels out the - sign.

jigglypuff314 (jigglypuff314):

don't forget parenthesis in that explanation ;)

OpenStudy (errinss):

THANK YOU PEOPLE!!!!!!!!! <3 i will be back tomorrow with more problems....that i need help with *SMILES* TANK CHU I WUV ALL OF CHU......<3

OpenStudy (cheez):

haha no prob.